A random variable has density function 2 2 / 3 for x 22 a. Find the expectation b. Variance
Answer #1

(M) f2) 2. 지그 2 3 o EX)= XfC에 일 } 2. 2 룷 da 3 2 L 2. 거2 = 그 [2] 2 2. + 2 E로 s x² fras) da. 21, 자 2 2[ log 2 =C -0.3010 = 1.39ucx)=f(x² - (fix)]2 - 1.3979-1 0.3979 variance = 0-3979